Select the equation of the line that passes through the given two points. Note that the correct equation may not be in slope-intercept form (solve for y to get the slope-intercept form).
12. (4,5), (19,-7)
A. -12x = -123 + 14y
B. -12x = -123 + 15y
C. -6x = 19 + 5y
D. 4x + 5y = 41
E. 5x + 5y = 10
F. -16x = 312 - 19y

Answer by Alan3354(69443)

Select the equation of the line that passes through the given two points. Note that the correct equation may not be in slope-intercept form (solve for y to get the slope-intercept form).
12. (4,5), (19,-7)
A. -12x = -123 + 14y
B. -12x = -123 + 15y
C. -6x = 19 + 5y
D. 4x + 5y = 41
E. 5x + 5y = 10
F. -16x = 312 - 19y
============================
You can sub for x & y in each equation, see which one fits.
---
Or, find the equation thru the 2 points then look for a match.